complete the ordered pairs for the equation 2x+t=10.
(5, ), ( , 10), ( ,-2), (0, )

Answer by checkley75(3666)

I'LL ASSUME THE T SHOULD BE Y.
SET X=5 2*5+Y=10 OR 1=+Y=10 OR Y=10-10 OR Y=0 THUS (5,0)
SET Y=10 2X+10=10
OR 2X=10-10 OR 2X=0 OR X=0(0,10)
SET Y=-2 2X-2=10 OR 2X=10=2 OR 2X=12 OR X=12/2 OR X=6 (6,-2)
SET X=0 2*0+Y=10 OR 0+Y=10 OR Y=10 (0,10)